OpenStudy (anonymous):

Evaluate the given function sin(A +B) by using the following information: sin A= 4/5 (1st Quad.) and cos B= -12/13 (2nd Quad.)

OpenStudy (anonymous):

sin(A+B) = sin A cos B + sin B cos A sin (A + B) = (4/5)(-12/13) + (5/13)(3/5)+ = -48/65 + 15/65 = -33/65

OpenStudy (anonymous):

Keep in mind that: If sin A = 4/5 then cos A = 3/5 just draw a right triangle, it is a 3-4-5 right triangle The other triangle is a 5-12-13 right triangle

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

Oh, Thanks so much!